# Calibrating LLM Confidence by Probing Perturbed Representation Stability

### v1

@@ I.1 Per-Dataset Aggregate Performance Tables
Meta-Llama-3.1-8B-Instruct P(True) 25.9$\pm$7.0 32.0$\pm$5.2 56.0$\pm$7.7 29.9$\pm$12.5 46.2$\pm$5.8
P(IK) 22.6$\pm$12.0 26.6$\pm$5.0 30.5$\pm$11.8 29.9$\pm$11.5 49.8$\pm$1.1
LTS 27.9$\pm$5.6 34.0$\pm$3.5 47.8$\pm$4.9 31.5$\pm$13.6 47.5$\pm$5.9
IT 27.8$\pm$5.9 33.2$\pm$4.6 53.9$\pm$6.1 30.6$\pm$13.3 47.2$\pm$6.0
SAPLMA-M 23.0$\pm$12.5 26.6$\pm$5.2 67.0$\pm$15.7 29.6$\pm$11.3 49.9$\pm$0.9
SAPLMA-UM 22.8$\pm$11.7 26.2$\pm$3.5 29.6$\pm$11.3 29.6$\pm$11.3 49.9$\pm$0.9
SAPLMA-F 22.5$\pm$11.5 26.1$\pm$3.5 29.7$\pm$11.4 29.7$\pm$11.4 49.8$\pm$1.4
CT 8.8$\pm$6.4 21.1$\pm$4.9 65.3$\pm$11.4 48.9$\pm$16.8 70.9$\pm$7.5
LitCab 8.8$\pm$7.6 22.5$\pm$4.8 65.3$\pm$9.1 46.2$\pm$13.8 66.0$\pm$9.6
CCPS 8.0$\pm$5.7 20.2$\pm$3.8 69.5$\pm$8.6 49.4$\pm$15.9 69.3$\pm$7.8
Qwen2.5-14B-Instruct P(True) 33.9$\pm$7.6 36.9$\pm$5.9 54.1$\pm$7.4 46.3$\pm$12.6 52.6$\pm$5.0
P(IK) 14.1$\pm$9.9 26.3$\pm$4.3 55.8$\pm$12.7 42.8$\pm$12.2 49.5$\pm$1.9
LTS 27.8$\pm$5.4 32.8$\pm$4.0 55.9$\pm$5.6 49.2$\pm$13.1 56.5$\pm$6.0
IT 33.6$\pm$6.1 36.7$\pm$4.8 55.0$\pm$6.0 47.5$\pm$13.9 54.0$\pm$5.9
SAPLMA-M 14.9$\pm$10.8 26.4$\pm$4.2 42.8$\pm$12.3 42.8$\pm$12.3 49.9$\pm$0.9
SAPLMA-UM 14.6$\pm$9.9 26.1$\pm$3.0 42.8$\pm$12.3 42.7$\pm$12.3 49.9$\pm$0.9
SAPLMA-F 14.7$\pm$10.1 26.2$\pm$3.3 42.8$\pm$12.3 42.8$\pm$12.4 49.9$\pm$1.0
CT 9.4$\pm$5.6 22.6$\pm$4.0 63.4$\pm$8.4 61.7$\pm$14.3 69.3$\pm$7.7
LitCab 34.4$\pm$10.3 37.0$\pm$7.3 49.4$\pm$10.1 56.8$\pm$13.4 62.5$\pm$6.8
CCPS 6.7$\pm$3.5 22.5$\pm$2.0 63.6$\pm$6.8 59.0$\pm$12.7 66.6$\pm$6.8
Mistral-Small-24B-Instruct-2501 P(True) 28.0$\pm$8.9 33.5$\pm$6.7 55.5$\pm$8.7 44.6$\pm$13.3 49.8$\pm$4.5
P(IK) 19.9$\pm$12.7 29.7$\pm$7.4 52.5$\pm$11.1 46.3$\pm$14.4 52.7$\pm$5.2
LTS 19.4$\pm$6.3 29.3$\pm$4.0 55.2$\pm$6.7 46.1$\pm$13.8 50.8$\pm$5.3
IT 26.2$\pm$7.9 32.5$\pm$5.6 55.2$\pm$7.4 45.5$\pm$13.6 50.6$\pm$4.5
SAPLMA-M 15.1$\pm$10.9 26.2$\pm$3.3 42.6$\pm$13.0 42.9$\pm$12.9 50.2$\pm$1.0
SAPLMA-UM 15.2$\pm$11.0 26.3$\pm$3.5 42.6$\pm$13.0 42.8$\pm$13.0 50.1$\pm$0.8
SAPLMA-F 14.9$\pm$10.9 26.2$\pm$3.4 42.6$\pm$13.0 42.7$\pm$13.0 50.0$\pm$1.6
CT 10.8$\pm$5.4 22.8$\pm$3.4 62.2$\pm$8.3 60.7$\pm$15.8 68.2$\pm$8.0
LitCab 11.2$\pm$5.0 24.6$\pm$3.1 60.2$\pm$6.8 60.5$\pm$13.3 66.4$\pm$6.5
CCPS 6.8$\pm$2.6 20.8$\pm$2.6 67.6$\pm$6.0 64.7$\pm$13.2 71.4$\pm$6.8
Qwen2.5-32B-Instruct P(True) 36.3$\pm$4.6 38.0$\pm$3.7 54.8$\pm$4.3 53.8$\pm$12.9 57.1$\pm$5.5
P(IK) 13.1$\pm$10.4 26.3$\pm$4.5 52.8$\pm$12.5 46.5$\pm$12.3 49.9$\pm$0.6
LTS 29.5$\pm$4.8 34.4$\pm$3.4 53.7$\pm$4.0 52.7$\pm$13.3 55.5$\pm$5.5
IT 33.2$\pm$7.1 37.3$\pm$5.3 52.7$\pm$6.8 49.0$\pm$12.7 51.6$\pm$4.5
SAPLMA-M 13.6$\pm$10.0 26.2$\pm$3.7 46.2$\pm$12.5 46.2$\pm$12.6 49.9$\pm$0.6
SAPLMA-UM 13.7$\pm$10.4 26.3$\pm$4.1 46.1$\pm$12.5 46.2$\pm$12.5 49.8$\pm$1.2
SAPLMA-F 13.7$\pm$10.2 26.3$\pm$3.8 46.1$\pm$12.5 46.2$\pm$12.5 49.8$\pm$0.8
CT 22.9$\pm$4.7 31.1$\pm$3.5 57.1$\pm$5.2 52.9$\pm$12.8 56.3$\pm$5.7
LitCab 28.4$\pm$8.1 33.2$\pm$5.5 52.7$\pm$8.5 60.2$\pm$13.0 62.3$\pm$7.6
CCPS 8.7$\pm$4.9 23.3$\pm$2.1 62.6$\pm$6.8 62.0$\pm$11.8 66.4$\pm$5.8
@@ I.3 Calibration Curves
I.4 Per-Task Performance Analysis
@@
I.5 Feature Importance Analysis
@@
To elucidate the contributions of various engineered features to the predictions of our CCPS model, we employed SHAP (SHapley Additive exPlanations) Lundberg and Lee (2017) (MIT License). This analysis utilized a model wrapper around our trained CCPS classifiers and a subset of the respective training data as background references for the shap.KernelExplainer with a logit link function. For Multiple-Choice (MC) models, which take a single feature vector as input, SHAP values directly indicate the importance of each of the $D_{f}$ features. The resulting "Feature-SHAP Correlation" plots (Figures 31 through 34 for MC model results) visualize the Pearson correlation between scaled feature values and their SHAP values, where colors typically distinguish positive and negative correlations, indicating how feature magnitudes influence the prediction towards correctness.

### v2

@@ Appendix I Evaluation on a High-Stakes Domain: MedMCQA
[…]
@@
J.4 Per-Task Performance Analysis
@@
Appendix K Ablation Studies
K.1 Token Importance in OE Models
To address the concern that using features from all tokens in an OE generation might be excessive, we conducted an ablation study to determine whether our convolutional architecture learns to prioritize semantically meaningful tokens over less informative "filler" words.
For this analysis, we used our final trained OE CCPS models. We performed a Token Masking Impact Analysis, which involved the following steps:
  1. 1.
    For each sample in our test set, we first obtained the baseline confidence score from the full, unmasked sequence of token features.
  2. 2.
    We then systematically masked one token at a time by zeroing out its entire feature vector and re-calculated the confidence score with the masked sequence.
  3. 3.
    The "impact" of each token was measured as the absolute difference between the baseline confidence and the confidence score from the sequence where that token was masked.
  4. 4.
    Finally, using Part-of-Speech (POS) tagging, we categorized each token as either a "Content" word (e.g., nouns, verbs, adjectives, adverbs) or a "Function" word (e.g., determiners, prepositions, pronouns) and compared the average impact scores for each category.
The results, summarized in Table 16, confirm that our convolutional architecture effectively learns to prioritize semantically meaningful tokens. Across all four LLMs, masking Content Words had a statistically significant and substantially larger impact on the final confidence score than masking Function Words. For example, for the Qwen2.5-14B and Qwen2.5-32B models, the impact of content words was over 75% greater than that of function words. This analysis demonstrates that the model’s sensitivity is not uniform across all tokens; the convolutional architecture effectively learns to place greater weight on semantically rich words while attenuating the influence of less informative ones, thus validating our approach to using the full feature sequence.
Base LLM Mean Impact of Content Mean Impact of Function Content vs. Function Impact Lift
Meta-Llama-3.1-8B 0.0182 0.0125 +45.6%
Qwen2.5-14B 0.0160 0.0091 +75.8%
Mistral-Small-24B 0.0385 0.0258 +49.2%
Qwen2.5-32B 0.0156 0.0088 +77.2%
Table 16: Impact of masking content vs. function words on the final confidence score. “Impact Lift” shows the percentage increase in impact when masking a content word compared to a function word.
K.2 Comparison with Self-Consistency Baseline
In response to reviewer feedback, and to provide a more comprehensive comparison against methods that rely on output sampling, we implemented and evaluated the Self-Consistency (SC) baseline Xiong et al. (2024). The SC method estimates confidence by generating multiple answer samples for a given question and using the consensus or frequency of the most common answer as the confidence score. We tested the SC method using our full experimental setup across all MMLU variants.
The results of this comparison are presented in Table 17. The findings show that while SC is a competitive baseline, CCPS consistently outperforms it, particularly in calibration metrics. As shown, CCPS achieves substantially lower ECE and Brier scores across all models and datasets, indicating significantly better calibration. Furthermore, CCPS generally demonstrates superior discriminative power, leading in AUCPR and AUROC in nearly all cases. These results further validate the effectiveness of our internal stability probing approach compared to methods based on external output consistency.
Table 17: Unified comparison of CCPS vs. Self-Consistency (SC) across three datasets: MMLU-CHOICE, MMLU-PRO-CHOICE, and MMLU-OE. Values are percentages. Best results per model are bolded.
Dataset Model Method ECE $\downarrow$ Brier $\downarrow$ ACC $\uparrow$ AUCPR $\uparrow$ AUROC $\uparrow$
MMLU-CHOICE Meta-Llama-3.1-8B SC 23.0 22.6 69.1 77.0 71.5
CCPS 6.5 17.1 73.4 84.1 77.1
Qwen2.5-14B SC 46.1 20.8 78.1 79.9 55.5
CCPS 6.3 13.1 80.2 92.1 81.6
Mistral-Small-24B SC 22.3 14.7 81.2 88.0 74.2
CCPS 5.8 11.5 83.0 93.1 83.3
Qwen2.5-32B SC 26.3 16.4 82.6 84.5 56.8
CCPS 6.3 10.8 84.1 94.1 82.8
MMLU-PRO-CHOICE Meta-Llama-3.1-8B SC 19.3 29.8 60.3 51.9 68.7
CCPS 4.5 20.0 70.4 55.2 67.9
Qwen2.5-14B SC 47.1 41.1 54.8 58.5 58.7
CCPS 4.2 20.1 69.2 75.8 74.0
Mistral-Small-24B SC 20.2 25.0 67.2 72.7 73.0
CCPS 4.5 18.6 71.3 79.5 77.2
Qwen2.5-32B SC 45.4 36.0 60.5 64.4 59.5
CCPS 4.6 18.5 71.8 82.4 77.8
MMLU-OE Meta-Llama-3.1-8B SC 11.2 21.2 71.2 46.6 67.0
CCPS 8.0 20.2 69.5 49.4 69.3
Qwen2.5-14B SC 22.5 33.0 56.6 51.7 59.8
CCPS 6.7 22.5 63.6 59.0 66.6
Mistral-Small-24B SC 14.0 24.1 65.8 58.9 67.2
CCPS 6.8 20.8 67.6 64.7 71.4
Qwen2.5-32B SC 23.7 33.6 56.0 54.3 59.8
CCPS 8.7 23.3 62.6 62.0 66.4
K.3 Disentangling Feature Contributions
To assess whether the observed performance gains of CCPS arise from its novel perturbation mechanism or merely from classification on features extracted from the LLM’s unperturbed representations, we conducted a comprehensive ablation study. This study aims to isolate and quantify the contribution of our perturbation-derived features.
While our main results demonstrate CCPS’s superior performance over other established probing baselines like SAPLMA, we designed this ablation to provide a more direct, controlled comparison within our own framework. We created and evaluated three variants of the CCPS model:
  • Original Only ($O$): A CCPS classifier trained exclusively on the 12 features derived from the LLM’s initial, unperturbed hidden state (e.g., original_entropy, original_log_prob_actual). This serves as our non-perturbation-based probe baseline.
  • Perturbation Only ($P$): A CCPS classifier trained exclusively on the 63 features derived from the perturbation process and its effects (e.g., epsilon_to_flip_token, pei_value_token, perturbed state statistics, and comparison features).
  • Full ($F$): Our complete CCPS model, which uses all 75 features (Original + Perturbation).
For each variant, the model architecture and training process were kept identical, with only the input dimension of the first layer adjusted to match the feature set size. The results for each LLM across the three MMLU variants are presented in Table 18. The results of this ablation study clearly demonstrate that our perturbation-based features are the primary driver of CCPS’s strong performance.
Dataset Model Variant ECE $\downarrow$ AUROC $\uparrow$
MMLU-CHOICE Meta-Llama-3.1-8B-Instruct $O$ 17.1 56.3
$P$ 11.3 63.0
$F$ 6.5 77.1
Qwen2.5-14B-Instruct $O$ 12.1 62.8
$P$ 8.6 74.7
$F$ 6.4 81.6
Mistral-Small-24B-Instruct-2501 $O$ 14.8 73.2
$P$ 8.1 76.5
$F$ 5.9 83.3
Qwen2.5-32B-Instruct $O$ 15.6 63.0
$P$ 8.9 77.8
$F$ 6.3 82.9
MMLU-PRO-CHOICE Meta-Llama-3.1-8B-Instruct $O$ 17.1 58.6
$P$ 6.1 61.2
$F$ 4.5 67.9
Qwen2.5-14B-Instruct $O$ 21.0 59.0
$P$ 11.0 70.7
$F$ 4.2 74.0
Mistral-Small-24B-Instruct-2501 $O$ 18.3 54.5
$P$ 8.9 63.1
$F$ 4.5 77.2
Qwen2.5-32B-Instruct $O$ 23.9 77.7
$P$ 6.5 77.5
$F$ 4.6 78.0
MMLU-OE Meta-Llama-3.1-8B-Instruct $O$ 14.6 48.7
$P$ 10.2 61.4
$F$ 8.0 69.3
Qwen2.5-14B-Instruct $O$ 15.2 60.1
$P$ 9.6 61.0
$F$ 6.8 66.6
Mistral-Small-24B-Instruct-2501 $O$ 10.8 50.1
$P$ 9.7 66.3
$F$ 6.8 71.4
Qwen2.5-32B-Instruct $O$ 23.4 52.2
$P$ 12.5 59.1
$F$ 8.7 66.4
Table 18: Ablation results of disentangling feature contributions. Variants are Original Only ($O$), Perturbation Only ($P$), and Full ($F$). Best performance per model is bolded.
Perturbation Features are the Dominant Performance Driver. As shown across all datasets, the $P$ model consistently and substantially outperforms the $O$ probe. This is particularly evident in ECE, where the perturbation-derived features consistently yield much better calibration than features from the unperturbed state. For example, on MMLU-PRO-CHOICE with Qwen2.5-14B, the $P$ model cuts the ECE in half (21.0% $\rightarrow$ 11.0%) and boosts AUROC by over 11 percentage points (59.0% $\rightarrow$ 70.7%) compared to the $O$ model. This directly isolates and confirms the significant contribution of our core perturbation mechanism.
The Value of Perturbation Increases with Task Difficulty. The performance gap between the perturbation-based models and the simple probe widens on more challenging datasets. While the $O$ probe is a reasonable baseline on standard MMLU-CHOICE, its performance degrades considerably on the more difficult MMLU-PRO-CHOICE and MMLU-OE tasks. In contrast, the $P$ model remains robust, showing much smaller performance degradation. For instance, with Meta-Llama-3.1-8B, the AUROC gap between $P$ and $O$ is 6.7 pp on MMLU-CHOICE, but this gap widens to 12.7 pp on the more challenging MMLU-OE. This strongly suggests that as task complexity increases and simple signals like initial log-probabilities become less reliable, the deeper stability signals captured by CCPS’s perturbation mechanism become critical for robust confidence estimation.
Features Combine Synergistically. Finally, the Full ($F$) model, which combines both feature sets, consistently achieves the best performance across all metrics and datasets. This indicates that the original state features, while less powerful on their own, provide complementary information that further refines the predictions made using the dominant perturbation-based features. This study provides clear evidence that the performance gains of CCPS are fundamentally driven by the novel perturbation mechanism we introduce, sharpening its distinction from prior probing methods.
K.4 Feature Importance Analysis with SHAP
@@
To investigate the contributions of various engineered features to the predictions of our CCPS model, we employed SHAP (SHapley Additive exPlanations) Lundberg and Lee (2017) (MIT License). This analysis utilized a model wrapper around our trained CCPS classifiers and a subset of the respective training data as background references for the shap.KernelExplainer with a logit link function. For Multiple-Choice (MC) models, which take a single feature vector as input, SHAP values directly indicate the importance of each of the $D_{f}$ features. The resulting "Feature-SHAP Correlation" plots (Figures 31 through 34 for MC model results) visualize the Pearson correlation between scaled feature values and their SHAP values, where colors typically distinguish positive and negative correlations, indicating how feature magnitudes influence the prediction towards correctness.
@@
K.5 Leave-One-Out Feature Ablation
To investigate the utility of our 75 features and address the possibility of redundancy, we performed a comprehensive leave-one-out feature ablation study. For this analysis, we retrained our CCPS model 75 times for each of the four base LLMs, each time with one feature removed, and evaluated the performance on both MMLU-CHOICE and the more challenging MMLU-PRO-CHOICE test sets. The importance of each feature was quantified by the resulting drop in AUROC when it was excluded from the model. Table 19 summarizes the five most and least influential features shared across all LLMs for each dataset, based on the average drop in AUROC.
Test Dataset Influential Features
MMLU-CHOICE Top 5 Most Influential:
perturbed_logit_argmax_mean
did_argmax_change_from_original_max
did_argmax_change_from_original_mean
pei_value_token
perturbed_logit_argmax_std
Top 5 Least Influential:
l2_dist_hidden_perturbed_from_original_std
cosine_sim_hidden_perturbed_to_original_std
original_norm_logits_L2
is_actual_token_original_argmax
jacobian_norm_token
MMLU-PRO-CHOICE Top 5 Most Influential:
js_div_perturbed_from_original_mean
cosine_sim_logits_perturbed_to_original_mean
js_div_perturbed_from_original_max
kl_div_perturbed_from_original_std
did_argmax_change_from_original_mean
Top 5 Least Influential:
perturbed_prob_actual_std
original_prob_actual
jacobian_norm_token
is_actual_token_original_argmax
original_entropy
Table 19: Top 5 most and least influential features identified for MMLU-CHOICE and MMLU-PRO-CHOICE. Each feature is shown on a separate line.
The Ablation Confirms that Features Measuring the Response to Perturbation are Most Important  Our analysis confirms the utility of our feature set, as removing any single feature did not improve performance across all model-task combinations, indicating a well-designed, non-redundant set. On MMLU-CHOICE, the most impactful features directly measure the outcome of the perturbation, such as whether the model’s top prediction changed (did_argmax_change...). This validates our core hypothesis that the stability of the LLM’s prediction under a targeted challenge is a key confidence signal.
The Nature of Informative Confidence Signals Shifts with Increasing Task Difficulty  Interestingly, when evaluating on the more challenging MMLU-PRO-CHOICE dataset, the nature of the most influential features shifts. The most vital shared features become those measuring the holistic distributional change under perturbation, such as js_div_perturbed_from_original_mean (the divergence between original and perturbed output distributions) and cosine_sim_logits_perturbed_to_original_mean (the similarity between original and perturbed logit vectors). This key insight suggests that on harder problems, quantifying the degree of representational shift appears to be a more robust indicator of confidence than tracking the stability of just the single top prediction.